# Testing the Star-disk Collision Model for Quasi-periodic Eruptions

### v1

@@ Abstract
Quasi-periodic eruptions (QPEs), the repeated outbursts observed in soft X-ray bands, have attracted broad interest, but their physical origin is under debate. One of the popular models, the star-disk collision model, suggests that QPEs can be produced through periodic collisions of an orbiting star with the accretion disk of a central black hole (BH). However, previous tests of the star-disk collision model mainly focus on the timing analysis. Other observed properties, such as peak luminosities $L_{\rm{p}}$, durations $t_{\rm{e}}$, and radiation temperatures $T_{\rm{p}}$ of the eruptions, are not systematically investigated. For a sample of six QPE sources and two QPE-like sources, we test the star-disk collision model by using these observables to derive the constraints on the stellar radius $R_{*}$. We find that, except for two sources (eRo-QPE3 and eRo-QPE4), the rest of the sample either has no allowed $R_{*}$ to simultaneously reproduce the observed $L_{\rm{p}}$ and $t_{\rm{e}}$, or the required $R_{*}$ is too large to avoid being disrupted by the central BH. For the two exceptions, a stellar radius of the order of $1\ R_{\rm{\odot}}$ is necessary to satisfy all the constraints. Another issue with the simplest version of this model is that it predicts $kT_{\rm{p}}\sim 10\ \rm{eV}$, one order of magnitude lower than the observed value.
@@ 1 Introduction
Recently, a new class of repeating soft X-ray outbursts called quasi-periodic eruptions (QPEs) have been discovered. Motivated by the first reported QPE, GSN 069 (Miniutti et al., 2019, 2023a, 2023b), RX J1301 and eRo-QPE1-4 are discovered (Sun et al., 2013; Giustini et al., 2020; Arcodia et al., 2021, 2024a, 2024b; Chakraborty et al., 2024). They exhibit luminous flares with the peak luminosity of $L_{\rm{p}}\approx 10^{42}\ \rm{erg/s}$, the duration of $t_{\rm{e}}\approx 1\ \rm{hr}$, and recurring over a period of $P\approx 10\ \rm{hr}$. Based on the fitting of the spectrum, their peak temperature $kT_{\rm{p}}\approx 100\ \rm{eV}$. They are located at the nuclei of their host galaxies with a quiescent luminosity of $L_{\rm{Q}}\approx\rm{10^{41}\ \rm{erg/s}}$, thus potentially linked to a massive BH with the mass of $M_{\rm{h}}\approx 10^{6}M_{\rm{\odot}}$. Also, due to periodic behaviors, QPE-like events like ASASSIN14-ko and Swift J0230 (hereafter called J0230 for short) (Payne et al., 2021, 2022, 2023; Huang et al., 2023; Linial & Quataert, 2024; Evans et al., 2023; Guolo et al., 2024) are expected to originate from the same physical process as QPEs.
@@
On the basis of observations, several theoretical models have been proposed to interpret QPEs. Generally, they can be classified into four kinds: 1) disk instability model (Raj et al., 2021; Pan et al., 2022; Kaur et al., 2023; Śniegowska et al., 2023), 2) self-lensing model in massive BH binary (Ingram et al., 2021), 3) star-disk collision model (Dai et al., 2010; Pihajoki, 2016; Suková et al., 2021; Xian et al., 2021; Franchini et al., 2023; Linial & Metzger, 2023; Linial & Quataert, 2024; Tagawa & Haiman, 2023), 4) mass transfer model between a stellar companion and the SMBH, such as via Roche lobe overflow (Metzger et al., 2022; Linial & Sari, 2023), repeated tidal stripping (King, 2020, 2022, 2023; Zhao et al., 2022; Chen et al., 2023; Wu et al., 2024), stream-stream collision induced by mass transfer (Krolik & Linial, 2022; Lu & Quataert, 2023). Here we focus on the star-disk collision model.
@@
For a star orbiting around the SMBH with an accretion disk, the star will keep colliding with a disk at the crossing points between the misaligned orbit and the disk plane. The flare comes out of the ejecta shocked by the collision. Dai et al. (2010) first investigated the star-disk collision model and built a connection between the quasi-periodic behavior of flares and the star’s orbital parameters. After the discovery of GSN 069 (Miniutti et al., 2019), the star-disk collision model was applied to interpret the QPEs. Several works have been dedicated to the timing analysis of QPE (Xian et al., 2021; Franchini et al., 2023; Linial & Metzger, 2023; Chakraborty et al., 2024; Zhou et al., 2024a, c, b; Miniutti et al., 2025), which aims to reproduce the high-low peak luminosity pattern and the long-short pattern of QPE by various precessions.
@@
Besides timing, the observed radiative property is also important for testing the model. Linial & Metzger (2023) studied the X-ray emission of the star-disk collision model by estimating $L_{\rm{p}}$, $T_{p}$, $t_{\rm{e}}$ and other observation properties. However, the star-disk collision model is still not fully constrained under the observation data. For example, it is difficult to find a satisfactory quiescent-level mass accretion rate to reproduce the observed $L_{\rm{p}}$, $T_{\rm{p}}$ and $t_{\rm{e}}$ simultaneously (Linial & Metzger, 2023). Several simulations try to figure out the details of the star-disk collision model (Vurm et al., 2024; Yao et al., 2025).
@@
In this paper, we carefully check the estimations in Linial & Metzger (2023) and try to obtain a more holistic constraint on the star-disk collision model based on various observations of QPEs and QPE-like events. §2 present a sample of QPE and QPE-like events to be used for the test. The theoretical framework of the star-disk collision model and the test method we apply are presented in §3, and the results are given in §4. Finally, we draw our conclusions in §5.
2 The Sample of QPE and QPE-like Events
@@
$\left(\rm{eV}\right)$
$\rm{event}$ $\rm{range}$ $\rm{typical\ value}$ $\rm{range}$ $\rm{typical\ value}$ $\rm{range}$ $\rm{typical\ value}$
[…]
@@
Table 1: The observed quantities regarding the radiation properties ($P$, $t_{\rm{e}}$, $L_{\rm{p}}$, $L_{\rm{Q}}$ and $kT_{\rm{p}}$) and the black hole masses $M_{\rm{h}}$ for a sample of QPEs and QPE-like events. The symbols in Ref. column are: a: Miniutti et al. (2019, 2023b, 2023a), b: Sun et al. (2013); Giustini et al. (2020), c: Arcodia et al. (2021); Chakraborty et al. (2024), d: Arcodia et al. (2021); Pasham et al. (2024b); Arcodia et al. (2024a), e: Arcodia et al. (2024b), f: Arcodia et al. (2024b), g: Payne et al. (2021, 2022, 2023); Huang et al. (2023); Linial & Quataert (2024), h: Evans et al. (2023); Guolo et al. (2024).
@@
QPE and QPE-like events are outstanding for their luminous flares with repeated pulse-like light curve profiles. In this section, we summarize these observed properties of QPE and QPE-like events, which are used in the constraints of $r_{*}$.
@@
GSN 069 shows a sequence of soft X-ray eruptions lasting for $t_{\rm{e}}\approx 1.1$ hr with $P\approx 8.5\ \rm{hr}$ and $L_{\rm{p}}\approx\left(10^{42}-10^{43}\right)\ \rm{erg/s}$ (Miniutti et al., 2019, 2023b, 2023a). Between eruptions, $L_{\rm{Q}}=\left(2-3\right)\times 10^{43}\ \rm{erg/s}$ (Miniutti et al., 2019, 2023b). Fitting by the Planck function, $kT_{\rm{p}}\approx\left(70-120\right)\ \rm{eV}$ (Miniutti et al., 2019, 2023b, 2023a). Based on the DISKBB model fitting, Miniutti et al. (2019, 2023b, 2023a) inferred $M_{\rm{h}}\approx\left(2-4\right)\times 10^{6}\rm{M}_{\rm{\odot}}$.
@@
RX J1301 has shown three rapid flares in $2019$ and one flare in $2000$. These flares last for $t_{\rm{e}}=\left(1-2\right)\ \rm{ks}$ with $L_{\rm{p}}\approx 10^{42}\ \rm{erg/s}$ and $P=17\ \rm{ks}$ (Giustini et al., 2020). During the quiescent state, $L_{\rm{Q}}=\left(0.3-2\right)\times 10^{41}\ \rm{erg/s}$.(Sun et al., 2013; Giustini et al., 2020) Based on diskbb or bbody models with serious components, $kT_{\rm{p}}=\left(94-330\right)\ \rm{eV}$ (Sun et al., 2013; Giustini et al., 2020). The mass of BH inferred from the UV/X-ray analysis is $M_{\rm{h}}=\left(8-28\right)\times 10^{5}\ M_{\rm{\odot}}$ (Sun et al., 2013).
@@
eRo-QPE1 is first reported for its luminous flares lasting for $t_{\rm{e}}=\left(6.6-8.6\right)\ \rm{hr}$ with $P=18\ \rm{hr}$ (Arcodia et al., 2021) and the later $3.5\ \rm{yr}$ monitoring presented by Chakraborty et al. (2024) shows its large variation in light curve. Generally, its various flares peak with $L_{\rm{p}}=\left(8-200\right)\times 10^{41}\ \rm{erg/s}$ and $L_{\rm{Q}}=\left(1-6.6\right)\times 10^{40}\ \rm{erg/s}$ (Arcodia et al., 2021; Chakraborty et al., 2024). eRo-QPE2 is also reported by Arcodia et al. (2021); Pasham et al. (2024b); Arcodia et al. (2024a) as a luminous QPE source with $P=2.4\ \rm{hr}$. It rises to a high state for $t_{\rm{e}}=\left(24-30\right)\ \rm{min}$ with $L_{\rm{p}}=\left(0.1-1.2\right)\times 10^{42}\ \rm{erg/s}$ (Arcodia et al., 2021; Pasham et al., 2024b; Arcodia et al., 2024a). In the quiescent state, $L_{\rm{Q}}=4\times 10^{40}\ \rm{erg/s}$ (Arcodia et al., 2021; Pasham et al., 2024b). For these two events, inferred by the zbbody, diskbb and tbabs $\times$ zbbody models, $kT_{\rm{p}}=\left(85-200\right)\ \rm{eV}$ for eRo-QPE1 (Arcodia et al., 2021; Chakraborty et al., 2024) and $kT_{\rm{p}}=\left(150-240\right)\ \rm{eV}$ for eRo-QPE2 (Arcodia et al., 2021; Pasham et al., 2024b; Arcodia et al., 2024a). Next, based on disk luminosity, duration, and period with adopting different viscosity parameters $\alpha$, $M_{\rm{h}}=\left(60-4\times 10^{6}\right)\ M_{\rm{\odot}}$ for eRo-QPE1 and $M_{\rm{h}}=\left(300-3\times 10^{6}\right)\ M_{\rm{\odot}}$ for eRo-QPE2 (Arcodia et al., 2021; Pasham et al., 2024b; Arcodia et al., 2024a).
@@
eRo-QPE3 and eRo-QPE4 are reported by Arcodia et al. (2024b). Based on observations, eRo-QPE3 peaks with $L_{\rm{p}}=\left(0.4-1\right)\times 10^{42}\ \rm{erg/s}$ with $P=20\ \rm{hr}$ and $L_{\rm{Q}}=1.6\times 10^{40}\ \rm{erg/s}$ (Arcodia et al., 2024b). For eRo-QPE4, the eruptions repeated every $14\ \rm{hr}$ with $L_{\rm{p}}=\left(4-15\right)\times 10^{42}\ \rm{erg/s}$ and $L_{\rm{Q}}=\left(10^{41}-10^{43}\right)\ \rm{erg/s}$ (Arcodia et al., 2024b). With assumption of a Gaussian light curve profile, $t_{\rm{e}}=\left(2-2.7\right)\ \rm{hr}$ for eRo-QPE3 and $t_{\rm{e}}=1\ \rm{hr}$ for eRo-QPE4 (Arcodia et al., 2024b). Also, fitting by a diskbb+pl model, $kT_{\rm{p}}=\left(70-140\right)\ \rm{eV}$ for eRo-QPE3 and $kT_{\rm{p}}=\left(79-127\right)\ \rm{eV}$ for eRo-QPE4 (Arcodia et al., 2024b). Moreover, inferring by the galaxy scaling relations, $M_{\rm{h}}=\left(4.6\times 10^{5}-4\times 10^{7}\right)\ M_{\rm{\odot}}$ for eRo-QPE3 and $M_{\rm{h}}=\left(1.3\times 10^{7}-1.2\times 10^{8}\right)\ M_{\rm{\odot}}$ for eRo-QPE4(Arcodia et al., 2024b).
@@
Similar to QPEs, ASASSN-14ko, belonging to a subset called repeated optical nuclear transients (RNOTs), is a famous QPE-like event. It shows repeated optical flares with $P\approx 115$ days in the nuclear of the galaxy (Payne et al., 2021). The flares last for $t_{\rm{e}}=\left(10-36\right)$ days with $L_{\rm{p}}\approx(0.3-10)\times 10^{44}\ \rm{erg/s}$ (Payne et al., 2021, 2022, 2023; Huang et al., 2023). Based on the relations between $M_{\rm{bulge}}-M_{\rm{BH}}$ and 2MASS photometry, Payne et al. (2021) obtained $M_{h}\approx 7\times 10^{7}\ \rm{M}_{\rm{\odot}}$. Using the blackbody model, $kT_{\rm{p}}=\left(0.3-2.5\right)\ \rm{eV}$ (Payne et al., 2021). The quiescent luminosity is $L_{\rm{Q}}=2\times 10^{42-43}\ \rm{erg/s}$ (Payne et al., 2021, 2022; Linial & Quataert, 2024).
@@
Furthermore, J0230 is another QPE-like source in the X-ray band with $P\approx 22\ \rm{days}$ (Evans et al., 2023; Guolo et al., 2024). Its period is much longer than that of QPE but shorter than that of RNOT. Based on observations. It peaks with $L_{\rm{p}}\approx 10^{42}\ \rm{erg/s}$ and $t_{\rm{e}}=\left(2.4-6.5\right)\ \rm{days}$ (Evans et al., 2023; Guolo et al., 2024) and its $L_{\rm{Q}}=\left(1-2\right)\times 10^{41}\ \rm{erg/s}$ (Evans et al., 2023). Except for the peculiar period of flares, its $kT_{\rm{p}}=150-200\ \rm{eV}$ and the mass of BH inferred from the $M_{\rm{h}}-\sigma_{*}$ scaling relation is about $\left(1.5\times 10^{6}-1\times 10^{7}\right)\ M_{\rm{\odot}}$ (Guolo et al., 2024).
@@
In Table 1, we collect several observable properties ($P$, $t_{\rm{e}}$, $L_{\rm{p}}$, $L_{\rm{Q}}$ and $kT_{\rm{p}}$) for a sample of QPEs and QPE-like events and their host galaxies’ central BH mass $M_{\rm{h}}$ estimated in the literature. For QPE sources that have been monitored for long-time coverage, the three most important observables ($t_{\rm{e}}$, $L_{\rm{p}}$ and $L_{\rm{Q}}$) usually vary from eruption to eruption, therefore their observed values each have a range. In order to simplify our analysis, we will use typical values for these observables.
3 Constraints for the Model

### v3

# Testing the Shock-cooling Emission Model from Star-disk Collisions for Quasi-periodic Eruptions

@@ Abstract
Quasi-periodic eruptions (QPEs), the repeated outbursts observed in soft X-ray bands, have attracted broad interest, but their physical origin is under debate. One of the popular models, the star-disk collision model, suggests that QPEs can be produced through periodic collisions of an orbiting star with the accretion disk of a central black hole (BH). However, previous tests of the star-disk collision model mainly focus on the timing analysis. Other observed properties, such as peak luminosities $L_{\rm{p}}$, durations $t_{\rm{e}}$, and radiation temperatures $T_{\rm{p}}$ of the eruptions, are not systematically investigated. For a sample of six QPE sources and two QPE-like sources, we test the shock-cooling emission model from star-disk collisions by using these observables to derive the constraints on the stellar radius $R_{\star}$. We find that, except for two sources (eRO-QPE3 and eRO-QPE4), the rest of the sample either has no allowed $R_{\star}$ to simultaneously reproduce the observed $L_{\rm{p}}$ and $t_{\rm{e}}$, or the required $R_{\star}$ is too large to avoid being disrupted by the central BH. For the two exceptions, a stellar radius of the order of $1\ R_{\rm{\odot}}$ is necessary to satisfy all the constraints. Another issue with the simplest version of this model is that it predicts $kT_{\rm{p}}\sim 10\ \rm{eV}$, one order of magnitude lower than the observed value.
@@ I Introduction
Recently, a new class of repeating soft X-ray outbursts called quasi-periodic eruptions (QPEs) have been discovered. Motivated by the first reported QPE, GSN 069 (Miniutti et al., 2019, 2023a, 2023b), RX J1301 and eRO-QPE1-4 are discovered (Sun et al., 2013; Giustini et al., 2020; Arcodia et al., 2021, 2024a, 2024b; Chakraborty et al., 2024). They exhibit luminous flares with the peak luminosity of $L_{\rm{p}}\approx 10^{42}\ \rm{erg/s}$, the duration of $t_{\rm{e}}\approx 1\ \rm{hr}$, and recurring over a period of $P\approx 10\ \rm{hr}$. Based on the fitting of the spectrum, their peak temperature $kT_{\rm{p}}\approx 100\ \rm{eV}$. They are located at the nuclei of their host galaxies with a quiescent luminosity of $L_{\rm{Q}}\approx\rm{10^{41}\ \rm{erg/s}}$, thus potentially linked to a massive BH with the mass of $M_{\rm{h}}\approx 10^{6}M_{\rm{\odot}}$. Also, due to periodic behaviors, QPE-like events like ASASSIN14-ko and Swift J0230 (hereafter called J0230 for short) (Payne et al., 2021, 2022, 2023; Huang et al., 2023; Linial and Quataert, 2024; Evans et al., 2023; Guolo et al., 2024) are expected to originate from the same physical process as QPEs.
@@
On the basis of observations, several theoretical models have been proposed to interpret QPEs. Generally, they can be classified into four kinds: 1) disk instability model (Raj et al., 2021; Pan et al., 2022; Kaur et al., 2023; Śniegowska et al., 2023), 2) self-lensing model in massive BH binary (Ingram et al., 2021), 3) star-disk collision model (Dai et al., 2010; Pihajoki, 2016; Suková et al., 2021; Xian et al., 2021; Franchini et al., 2023; Lu and Quataert, 2023; Linial and Metzger, 2023; Linial and Quataert, 2024; Tagawa and Haiman, 2023; Linial et al., 2025), 4) mass transfer model between a stellar companion and the SMBH, such as via Roche lobe overflow (Metzger et al., 2022; Linial and Sari, 2023), repeated tidal stripping (King, 2020, 2022, 2023; Zhao et al., 2022; Chen et al., 2023; Wu et al., 2024), stream-stream collision induced by mass transfer (Krolik and Linial, 2022). Here we focus on the star-disk collision model.
@@
For a star orbiting around the SMBH with an accretion disk, the star will keep colliding with a disk at the crossing points between the misaligned orbit and the disk plane. The flare comes out of the ejecta shocked by the collision. Dai et al. (2010) first investigated the star-disk collision model and built a connection between the quasi-periodic behavior of flares and the star’s orbital parameters. After the discovery of GSN 069 (Miniutti et al., 2019), the star-disk collision model was applied to interpret the QPEs. Several works have been dedicated to the timing analysis of QPE (Xian et al., 2021; Franchini et al., 2023; Linial and Metzger, 2023; Chakraborty et al., 2024; Zhou et al., 2024a, c, b; Miniutti et al., 2025), which aims to reproduce the high-low peak luminosity pattern and the long-short pattern of QPE by various precessions.
Besides timing, the observed radiative property is also important for testing the model. By estimating $L_{\rm{p}}$, $T_{p}$, $t_{\rm{e}}$ and other observation properties, Linial and Metzger (2023) analytically studied the X-ray emission of the star-disk collision model. Tagawa and Haiman (2023) investigates the break-out emission in the star-disk collisions model, incorporating the effect of orbital inclination. Vurm et al. (2024) explored the photon production process during the star-disk collisions and the radiative properties of the shocked material when it freely expands from the disk surface.
As an improved version fo the star-disk collision model, the stream-disk collision scenario was studied by Lu and Quataert (2023), Linial and Quataert (2024) and Linial et al. (2025). Here, the debris stream was shed from the star either due to heating-induced inflation and ablation resulted from a previous collision, as shown in the hydrodynamic simulation by Yao et al. (2025), or because of the tidal stripping by the SMBH (Huang et al., 2023).
@@
However, the star-disk collision model is still not fully constrained under the observation data, much less the validation of its improved version : the stream-disk collision scenario. For example, it is difficult to find a satisfactory quiescent-level mass accretion rate to reproduce the observed $L_{\rm{p}}$, $T_{\rm{p}}$ and $t_{\rm{e}}$ simultaneously, as was shown in Linial and Metzger (2023).
@@
In this paper, we carefully check the estimations in Linial and Metzger (2023) and try to obtain a more holistic constraint on the shock-cooling emission model from star-disk collisions based on various observations of QPEs and QPE-like events. §II present a sample of QPE and QPE-like events to be used for the test. The theoretical framework of the model and the test method we apply are presented in §III, and the results are given in §IV. Some additional effects and possible scenarios are discussed in $\S$ V. Finally, we draw our conclusions in §VI.
II The Sample of QPE and QPE-like Events
@@
$\rm{Source}$ $\rm{range}$ $\rm{typical\ value}$ $\rm{range}$ $\rm{typical\ value}$ $\rm{range}$ $\rm{typical\ value}$ $\left(\rm{eV}\right)$
$\rm{GSN}\ \rm{069}^{a}$ $8.6\ \rm{hr}$ 11-55 33 $(0.7$-$1.2)\ \rm{hr}$ $0.94\ \rm{hr}$ 200200-$300$ $250$ 7070-$120$ 0.30.3-44
$\rm{RX}\ \rm{J1301}^{b}$ $4.7\ \rm{hr}$ 11-33 $1.0$ $(0.3$-$0.5)\ \rm{hr}$ $0.4\ \rm{hr}$ 0.30.3-22 $0.6$ 9494-$330$ 0.80.8-$2.8$
$\rm{eRO}$-$\rm{QPE1}^{c}$ $18\ \rm{hr}$ 0.370.37-$43$ $4.0$ $(6.6$-$8.6)\ \rm{hr}$ $7.5\ \rm{hr}$ 0.100.10-$0.66$ $0.38$ 8585-$200$ $6\times 10^{-5}$-44
$\rm{eRO}$-$\rm{QPE2}^{d}$ $2.4\ \rm{hr}$ 0.10.1-$1.2$ $0.35$ $(0.4$-$0.5)\ \rm{hr}$ $5\ \rm{hr} - 0.40.4$ 150150-$240$ $\times 10^{-5}-33$
$\rm{eRO}$-$\rm{QPE3}^{e}$ $20\ \rm{hr}$ 0.40.4-11 $0.6$ $(2$-$2.7)\ \rm{hr}$ $2.3\ \rm{hr}$ 0.20.2-$30$ $2.5$ 7070-$140$ 0.460.46-$40$
$\rm{eRO}$-$\rm{QPE4}^{f}$ $14\ \rm{hr}$ 44-$15$ $7.9$ $(0.9$-$1.1)\ \rm{hr}$ $\rm{hr} 11-$$100$ $10$ 7979-$127$ 1313-$120$
$\rm{ASASSN}$-$\rm{14ko}^{g}$ $115\ \rm{days}$ 1212-$200$ $50$ $(10$-$36)\ \rm{days}$ $23\ \rm{days}$ 2020-$200$ $63$ 0.30.3-$2.5$ 1212-$159$
$\rm{J0230}^{h}$ $22\ \rm{days}$ 0.70.7-66 22 $(2.4$-$6.5)\ \rm{days}$ $4\ \rm{days}$ 11-22 $1.5$ 150150-$200$ 1.51.5-$10$
@@
Table 1: The observed quantities regarding the radiation properties ($P$, $t_{\rm{e}}$, $L_{\rm{p}}$, $L_{\rm{Q}}$ and $kT_{\rm{p}}$) and the black hole masses $M_{\rm{h}}$ for a sample of QPEs and QPE-like events. The symbols in Ref. column are: a: Miniutti et al. (2019, 2023b, 2023a), b: Sun et al. (2013); Giustini et al. (2020), c: Arcodia et al. (2021); Chakraborty et al. (2024), d: Arcodia et al. (2021); Pasham et al. (2024); Arcodia et al. (2024a), e: Arcodia et al. (2024b), f: Arcodia et al. (2024b), g: Payne et al. (2021, 2022, 2023); Huang et al. (2023); Linial and Quataert (2024), h: Evans et al. (2023); Guolo et al. (2024).
@@
QPE and QPE-like events are outstanding for their luminous flares with repeated pulse-like light curve profiles. In this section, we summarize these observed properties of QPE and QPE-like events, which are used in the constraints of $r_{\star}$.
@@
GSN 069 shows a sequence of soft X-ray eruptions lasting for $t_{\rm{e}}\approx 1.1$ hr with $P\approx 8.5\ \rm{hr}$ and $L_{\rm{p}}\approx\left(10^{42}-10^{43}\right)\ \rm{erg/s}$ (Miniutti et al., 2019, 2023b, 2023a). Between eruptions, $L_{\rm{Q}}=\left(2-3\right)\times 10^{43}\ \rm{erg/s}$. Spectral fitting gives $kT_{\rm{p}}\approx\left(70-120\right)\ \rm{eV}$. Based on the DISKBB model fitting, Miniutti et al. (2019, 2023b, 2023a) inferred $M_{\rm{h}}\approx\left(2-4\right)\times 10^{6}\rm{M}_{\rm{\odot}}$.
@@
RX J1301 has shown three rapid flares in $2019$ and one flare in $2000$. These flares last for $t_{\rm{e}}=\left(1-2\right)\ \rm{ks}$ with $L_{\rm{p}}\approx 10^{42}\ \rm{erg/s}$ and $P=17\ \rm{ks}$ (Giustini et al., 2020). During the quiescent state, $L_{\rm{Q}}=\left(0.3-2\right)\times 10^{41}\ \rm{erg/s}$ (Sun et al., 2013; Giustini et al., 2020). The mass of BH inferred from the UV/X-ray analysis is $M_{\rm{h}}=\left(8-28\right)\times 10^{5}\ M_{\rm{\odot}}$ (Sun et al., 2013).
@@
eRO-QPE1 is first reported for its luminous flares lasting for $t_{\rm{e}}=\left(6.6-8.6\right)\ \rm{hr}$ with $P=18\ \rm{hr}$ (Arcodia et al., 2021) and the later $3.5\ \rm{yr}$ monitoring presented by Chakraborty et al. (2024) shows its large variation in light curve. Generally, its various flares peak with $L_{\rm{p}}=\left(8-200\right)\times 10^{41}\ \rm{erg/s}$ and $L_{\rm{Q}}=\left(1-6.6\right)\times 10^{40}\ \rm{erg/s}$. eRO-QPE2 is reported by Arcodia et al. (2021); Pasham et al. (2024); Arcodia et al. (2024a) as a luminous QPE source with $P=2.4\ \rm{hr}$. It rises to a high state for $t_{\rm{e}}=\left(24-30\right)\ \rm{min}$ with $L_{\rm{p}}=\left(0.1-1.2\right)\times 10^{42}\ \rm{erg/s}$. In the quiescent state, $L_{\rm{Q}}=4\times 10^{40}\ \rm{erg/s}$. The central BH masses were estimated for these two sources based on disk luminosity, duration, and period with adopting different viscosity parameters $\alpha$.
@@
eRO-QPE3 and eRO-QPE4 are reported by Arcodia et al. (2024b). Based on observations, eRO-QPE3 peaks with $L_{\rm{p}}=\left(0.4-1\right)\times 10^{42}\ \rm{erg/s}$ with $P=20\ \rm{hr}$ and $L_{\rm{Q}}=1.6\times 10^{40}\ \rm{erg/s}$. For eRO-QPE4, the eruptions repeated every $14\ \rm{hr}$ with $L_{\rm{p}}=\left(4-15\right)\times 10^{42}\ \rm{erg/s}$ and $L_{\rm{Q}}=\left(10^{41}-10^{43}\right)\ \rm{erg/s}$. Moreover, their central BH masses are inferred by the galaxy scaling relations (Arcodia et al., 2024b).
@@
Similar to QPEs, ASASSN-14ko, belonging to a subset called repeated optical nuclear transients (RNOTs), is a famous QPE-like event. It shows repeated optical flares with $P\approx 115$ days in the nuclear of the galaxy (Payne et al., 2021). The flares last for $t_{\rm{e}}=\left(10-36\right)$ days with $L_{\rm{p}}\approx(0.3-10)\times 10^{44}\ \rm{erg/s}$ (Payne et al., 2021, 2022, 2023; Huang et al., 2023). Spectral fitting gives $kT_{\rm{p}}=\left(0.3-2.5\right)\ \rm{eV}$ and the quiescent luminosity is $L_{\rm{Q}}=2\times 10^{42-43}\ \rm{erg/s}$. Based on the relations between $M_{\rm{bulge}}-M_{\rm{BH}}$ and 2MASS photometry, Payne et al. (2021) obtained $M_{h}\approx 7\times 10^{7}\ \rm{M}_{\rm{\odot}}$.
@@
Furthermore, J0230 is another QPE-like source in the X-ray band with $P\approx 22\ \rm{days}$ (Evans et al., 2023; Guolo et al., 2024). Its period is much longer than that of QPE but shorter than that of RNOT. It peaks with $L_{\rm{p}}\approx 10^{42}\ \rm{erg/s}$ and $t_{\rm{e}}=\left(2.4-6.5\right)\ \rm{days}$ and its $L_{\rm{Q}}=\left(1-2\right)\times 10^{41}\ \rm{erg/s}$. Except for the peculiar period of flares, its $kT_{\rm{p}}=150-200\ \rm{eV}$ and the mass of BH inferred from the $M_{\rm{h}}-\sigma_{*}$ scaling relation is about $\left(1.5\times 10^{6}-1\times 10^{7}\right)\ M_{\rm{\odot}}$ (Guolo et al., 2024).
@@
In Table 1, we collect several observable properties ($P$, $t_{\rm{e}}$, $L_{\rm{p}}$, $L_{\rm{Q}}$ and $kT_{\rm{p}}$) for a sample of QPEs and QPE-like events and their host galaxies’ central BH mass $M_{\rm{h}}$ estimated in the literature. For QPE sources that have been monitored for long-time coverage, the three most important observables ($t_{\rm{e}}$, $L_{\rm{p}}$ and $L_{\rm{Q}}$) usually vary from eruption to eruption, therefore their observed values each have a range. In order to simplify our analysis in $\S$ IV, we will use typical values for these observables.
III Constraints for the Model